OpenStudy (anonymous):

A graphing calculator is not allowed for the following problems. ________________________________________ 1.Write an equation for the line tangent to the curve x2 + 4x + y3 = 4 at the point (1, -1).

OpenStudy (anonymous):

y'=((-2/3)x-1)(-x^2-4x+4)^-2/3..............chain rule and stuff jus took the first derivative

OpenStudy (anonymous):

implcit diff for this one

OpenStudy (anonymous):

\[x^2 + 4x + y^3 = 4 \] \[2x+4+3y^2y'=0\] \[y'=\frac{2x+4}{3y^2}\] replace x by 1, y by -1 to find your slope, then use the point - slope formula